If the graph of the parabola y=kx^2+6x+k+8 is tangent to x-axis, then what is the positive value of k?: If the graph of the parabola y=kx^2+6x+k+8 is tangent to x-axis, then what is the positive value of k?
Answer by oscargut(682)

if the graph is tangent to x-axis then discriminant is 0
then b^2-4ac=0
then
6^2-4(k)(k+8)=0
36-4k^2-32k=0
-4k^2-32k+36=0
k^2+8x-9=0
then k=1 or k=-9
then k=1 (positive value)
Answer: k=1